OpenStudy (anonymous):

hi. i dealing with order of operations agreement my problem is 6+ top of the bar is 16-4 bleow is 2x2+2 and -2

OpenStudy (valpey):

Is this what you mean? \[6+\frac{16-4}{2x2+2}-2\]

OpenStudy (anonymous):

YES

OpenStudy (valpey):

Step 1: 2*2 = 4 Step 2: treat top and bottom of fraction like they are in parentheses Step 3: simplify fraction Step 4: add across

OpenStudy (valpey):

\[6+\frac{16-4}{2*2+2}-2\] \[6+\frac{16-4}{4+2}-2\] \[6+\frac{12}{6}-2\] \[6+2-2\] \[6\]

OpenStudy (anonymous):

Thanks i see where i got stuck at. am i able to wirte my problem that way you did on here?

OpenStudy (valpey):

Yes. You can use the equation editor and use "\frac{top}{bottom}" to create long fractions

OpenStudy (anonymous):

ok thanks this is my first time using this service

OpenStudy (valpey):

use backslash openbracket: \[ at the beginning of an expression then backslash closedbracket \ ] at the end (without the space)